OpenStudy (anonymous):

Part 1: Determine whether 2 is a zero of the polynomial P(x) = 2x3 – 5x2 + 4x – 6 by using the Remainder Theorem. Show your work. (4 points) Part 2: Explain how the Remainder Theorem is useful in finding the zeros of a polynomial function. (4 points)

OpenStudy (anonymous):

solve for p(2) for the part one, if p(2)=0 then, 2 is a factor but if it is not equal to zero, then 2 is not a factor(the output will be the remainder).

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\(P(2)=2(2)^3-5(2)^2+4(2)-6=\)

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

brb.. need to eat my dinner XD

OpenStudy (anonymous):

p(2)=2(4) -5 (2) +4 (2) -6 p(2)= 8 -10 +8 -6 p(2)= 0

OpenStudy (loser66):

that means 2 is one of zero. share me 2 points hehehe

OpenStudy (anonymous):

ok